### v1

# Simplifying generic smooth maps to the $2$–sphere and to the plane

Abstract.
We study how to construct explicit deformations of generic smooth maps from closed $n$–dimensional manifolds $M$ with $n\geq 4$ to the $2$–sphere $S^{2}$ and show that every smooth map $M\to S^{2}$ is homotopic to a $C^{\infty}$ stable map with at most one cusp point and with only folds of the middle absolute index. Furthermore, if $n$ is even, such a $C^{\infty}$ stable map can be so constructed that the restriction to the singular point set is a topological embedding. As a corollary, we show that for $n\geq 4$ even, there always exists a $C^{\infty}$ stable map $M\to\mathbb{R}^{2}$ with at most one cusp point such that the restriction to the singular point set is a topological embedding. As another corollary, we give a new proof to the existence of an open book structure on odd dimensional manifolds which extends a given one on the boundary, originally due to Quinn. Finally, using the open book structure thus constructed, we show that $k$–connected $n$–dimensional manifolds always admit a fold map into $\mathbb{R}^{2}$ without folds of absolute indices $i$ with $1\leq i\leq k$, for $n\geq 7$ odd and $21\leq k<(n-3)/$ 2.
1. Introduction
For an arbitrary smooth closed manifold $M$ of dimension $n\geq 1$, we have a Morse function $f:M\to\mathbb{R}$. It is known that by modifying $f$ by homotopy, we can arrange so that the critical values are in the order of indices, i.e. for every pair $p,q$ of critical points of indices $i(p),i(q)$, respectively, with $i(p)<i(q)$, we have $f(p)<f(q)$. Furthermore, Smale [20] showed that if an $n$–dimensional closed manifold $M$ with $n\geq 6$ is $k$–connected with $k\geq 1$, then there exists a Morse function $f:M\to\mathbb{R}$ without critical points of indices $i$ with $1\leq i\leq k$ and $n-k\leq i\leq n-1$.
So, the following problem naturally arises.
Problem 1.1.
How about generic maps (or $C^{\infty}$ stable maps) to manifolds of dimensions $\geq 2$? The singular values can be nicely arranged with respect to the indices? If the manifold is highly connected, then the singular points of certain indices can be eliminated?
We note that such a problem has been considered in [5], for example. Recall that for a $C^{\infty}$ stable map $f$ of an $n$–dimensional manifold $M$, $n\geq 2$, into a manifold of dimension $2$, we have only folds and cusps as its singularities, and each of them has an absolute index (for details, see §2), which is a concept similar to that for a nondegenerate critical point of a smooth function. Furthermore, the singular point set $S(f)$ is a closed $1$–dimensional submanifold of $M$.
In the present paper, we consider the above problem for generic smooth maps to $S^{2}$ and to $\mathbb{R}^{2}$. Our first result is Theorem 4.3: for a closed connected $n$–dimensional manifold $M$ with $n\geq 3$, there always exists a $C^{\infty}$ stable map $f:M\to S^{2}$ such that for $n$ odd, $f$ has only folds of absolute index $(n-1)/2$ and has no cusps and that for $n$ even, $f$ has only folds of absolute index $(n-2)/2$ and at most one cusp, whose absolute index is equal to $(n-2)/2$. Furthermore, when $n$ is even, we can choose $f$ such that the restriction of $f$ to the singular point set $S(f)$ is a topological embedding. In the case of $n$ even, using the latter result, we will show, as Theorem 5.1, that there always exists a $C^{\infty}$ stable map $f:M\to\mathbb{R}^{2}$ with at most one cusp point such that $f|_{S(f)}$ is a topological embedding.
In the case of $n$ odd, we will also show that we can construct a $C^{\infty}$ stable map $f:M\to\mathbb{R}^{2}$ without cusps such that the singular value set winds around the origin. This leads to a new proof to the existence of an open book structure on odd dimensional smooth manifolds (see Theorem 6.3), which is originally due to Winkelnkemper [25], Tamura [22], Lawson [10] and Quinn [14]. In fact, we will show that any open book structure on the boundary $\partial M$ of a compact odd dimensional manifold $M$ extends through $M$. The open book that we construct has a very nice property with respect to radial Morse function, and this observation will be used to prove Theorems 7.2 and 7.6: if $M$ is a closed $k$–connected $n$–dimensional manifold with $n\geq 7$ odd and $21\leq k\leq(n-5)/$ 2, then there exists a $C^{\infty}$ stable map $f:M\to\mathbb{R}^{2}$ without cusp points such that $f$ has no fold points of absolute index $i$ with $1\leq i\leq k$. This gives a good answer to the second half of Problem 1.1.
The paper is organized as follows. In §2, we review important properties of generic maps of manifolds of dimension $n\geq 2$ into $2$–dimensional manifolds. In particular, we give some lists of changes of the singular value set for a generic $1$–parameter family of such maps, which will play an essential role in this paper. In fact, in the case of manifolds $M$ of dimension $n=4$, a similar set of lists has been given in [2] (see also [1]). In §3, we introduce the notion of moves, which change the singular value set of a generic smooth map. We rigorously formulate the notion of an always-realizable move, which was originally introduced in [2], as a formal change of the singular value set that is always realized by a generic $1$–parameter family of smooth maps.
In §4, using such always-realizable moves, we try to arrange the singular value set of generic maps into $S^{2}$. We first arrange them so that they wind around a certain point in the target, and then we arrange them in the order of the absolute index. These explicit procedures using always-realizable moves lead to the proof of Theorem 4.3. In §5, using such a generic smooth map into $S^{2}$, in the even dimensional case, we construct a generic smooth map into $\mathbb{R}^{2}$ and prove Theorem 5.1. The idea of the proof of these theorems are basically similar to that exploited in dimension $4$ in [2]. However, in dimension $4$, one can use Lefschetz singularities (see Remark 4.12). In our case of arbitrary dimensions, we will give new techniques that do not involve Lefschetz singularities. This is one of the new contributions of this paper.
In §6, for a compact manifold $M$ of an odd dimension with an open book structure being given on the boundary, we construct a smooth generic map into $\mathbb{R}^{2}$ which is compatible to the open book structure on the boundary, and, using techniques employed in §4, we arrange the singular value set so that it winds around the origin. Then, we prove the existence of an open book structure on $M$ which extends the given one on $\partial M$. In §7, we study the connectivity of the page and the binding of the open book structure thus constructed, and using the Cerf theory techniques developed in [8], we eliminate fold loci of certain absolute indices.
Throughout the paper, all manifolds and maps are smooth of class $C^{\infty}$ unless otherwise specified. For a smooth map $f:M\to N$ between smooth manifolds, $S(f)(\subset M)$ denotes the set of singular points of $f$. For a real number $x$, $\lfloor x\rfloor$ denotes the greatest integer not exceeding $x$.
2. Generic maps into surfaces
Let $M$ be a smooth closed manifold of dimension $n\geq 2$, and $N$ a smooth surface.
Definition 2.1.
A singular point $p$ of a smooth map $f:M\to N$ that can be described by the normal form
$(t,x_{1},x_{2},\ldots,x_{n-1})\mapsto(X,Y)=(t,-x_{1}^{2}-\cdots-x_{i}^{2}+x_{i+1}^{2}+\cdots+x_{n-1}^{2})$
with respect to appropriate coordinates around $p$ and $f(p)$ is called a fold singularity (or a fold), where the integer $i$ is called the index with respect to the $(-Y)$–direction. When we choose the $Y$–coordinate so that $0\leq i\leq\left\lfloor(n-1)/2\right\rfloor$ holds, $i$ is called the absolute index. [^1] We warn the reader that this convention is different from that adopted in [11, p. 273], [26], etc.
When $i=0$ or $n-1$, the singular point is called a definite fold. Otherwise, it is called an indefinite fold.
Definition 2.2.
A singular point of a smooth map $M\to N$ that has the normal form
$(t_{1},t_{2},x_{1},x_{2},\ldots,x_{n-2})$
$\mapsto(X,Y)=(t_{1},t_{2}^{3}+t_{1}t_{2}-x_{1}^{2}-\cdots-x_{i}^{2}+x_{i+1}^{2}+\cdots+x_{n-2}^{2})$
is called a cusp singularity (or a cusp), where $i$ is called the index with respect to the $(-Y)$–direction. When we choose the $Y$–coordinate so that $0\leq i\leq\left\lfloor(n-2)/2\right\rfloor$ holds, $i$ is called the absolute index.
A smooth map $f:M\to N$ that has only folds and cusps as its singularities is called a generic map (or an excellent map). On the other hand, if $f$ has only fold singularities, then it is called a fold map. [^2] In [6], this is called a submersion with folds. It is well known that a smooth map $M\to N$ can always be approximated by a generic map arbitrarily well and hence is homotopic to a generic map [23, 24].
For a smooth map $f:M\to N$, $S(f)$ denotes the set of singular points of $f$. If $f$ is generic and the source manifold $M$ is closed (compact and without boundary), then $S(f)$ is a disjoint union of finitely many circles, consisting of finitely many cusps together with finitely many arcs and circles of folds, and $f|_{S(f)\setminus\{\mbox{\footnotesize{cusps}}\}}$ is an immersion.
In the following, we assume that the source manifold is closed unless otherwise stated.
Definition 2.3.
A generic map $f:M\to N$ of an $n$–dimensional manifold $M$ with $n\geq 2$ into a surface $N$ is a $C^{\infty}$ stable map (or a stable map, for short) if the following conditions are satisfied.
  1. (1)
    The restriction $f|_{S(f)\setminus\{\mbox{\footnotesize{cusps}}\}}$ is an immersion with normal crossings, and
  2. (2)
    for every cusp $p\in M$, we have $f^{-1}(f(p))\cap S(f)=\{p\}$.
Note that $C^{\infty}$ stable maps are usually defined in another way (see [6]): however, it is known that the above definition is equivalent to such a usual definition (see, for example, [12]). Note also that the set of stable maps is open and dense in the space $C^{\infty}(M,N)$ of all smooth maps endowed with the Whitney $C^{\infty}$–topology.
Let $f:M\to N$ be a generic map. For a fold $p\in M$ of index $i$ with respect to the $(-Y)$–direction in the sense of Definition 2.1, the immersion $f|_{S(f)}$ near $p$ is normally oriented so that $f|_{f^{-1}(\ell)}:f^{-1}(\ell)\to\ell$ has $p$ as a non-degenerate critical point of index $i$, where $\ell$ is a small open arc embedded in $N$ which intersects $f(S(f))$ transversely at exactly one point $f(p)$, and the normal orientation corresponds to the negative direction of $\ell\cong\mathbb{R}$, i.e. the $(-Y)$–direction (see Fig. 1).
[…]
Figure 1. Normal orientation for $f|_{S(f)}$, depicted as thick downward arrow
Note that if we reverse the normal orientation, then the index $i$ changes to $(n-1)-i$. Therefore, if we require that $i$ should be the absolute index, then the normal orientation is uniquely determined, except for the case where $n$ is odd and $i=(n-1)/2$. When the absolute index satisfies $i=(n-1)/2$, both of the normal orientations give the same index. In this case, we usually do not associate any normal orientation (or, in other words, both normal orientations are associated, or the normal orientation is redundant).
For the image of a cusp point, we can also consider a “normal orientation”: this is considered to be normal to the two branches of curves emanating from the point. The indices around the image of a cusp are always as depicted in Fig. 2 for some integer $i$ with respect to appropriate normal orientations as in the figure.
[…]
Figure 2. Normal orientations and indices around the image of a cusp
We see easily that the absolute indices around cusps and the corresponding normal orientations are as indicated in Fig. 3. Note that when $n$ is odd, Case (2) can occur, while when $n$ is even, Case (3) can occur. Here, we usually do not indicate the normal orientation of the image of a cusp. This is because for Case (1), the normal orientation is considered to be the unique one consistent with the normal orientations for the two branches of curves (as in Fig. 2), for Case (2), it should be considered to be the unique one consistent with the normal orientation of one of the two branches, and for Case (3), both of the normal orientations can be chosen.
[…]
Figure 3. Absolute indices and normal orientations around cusp images
3. Diagrams and their moves
In this section, we introduce the notion of diagrams and moves for them. In the following, we assume that the source manifold $M$ is closed of dimension $n\geq 2$ and that $N$ is a smooth surface.
Definition 3.1.
Let $S$ be a finite disjoint union of circles and $\gamma:S\to N$ a smooth map with the following properties.
  1. (1)
    The map $\gamma$ is an immersion with normal crossings outside of a finite set $C(\subset S)$.
  2. (2)
    Around each point $c_{\ell}$ of $C$, $\gamma$ is of the form $t\mapsto(t^{2},t^{3})$ with respect to appropriate local coordinates around $c_{\ell}$ and $\gamma(c_{\ell})$.
  3. (3)
    For each point $c_{\ell}\in C$, we have $\gamma^{-1}(\gamma(c_{\ell}))=\{c_{\ell}\}$.
We also assume that the immersion obtained by restricting $\gamma$ to each component $S_{k}$ of $S\setminus C$ is normally oriented. Moreover, to each $S_{k}\subset S\setminus C$ is associated an integer $i_{k}$ with $0\leq i_{k}\leq\lfloor(n-1)/2\rfloor$, and to each $c_{\ell}\in C$ is associated an integer $j_{\ell}$ with $0\leq j_{\ell}\leq\lfloor(n-2)/2\rfloor$ in such a way that around each point of $C$, they are as depicted in Fig. 3. When $n$ is odd and $S_{k}$ has index $i_{k}=(n-1)/2$, it is allowed that no normal orientation is given for $\gamma|_{S_{k}}$. Then, we denote by $\mathcal{D}$ the image $\gamma(S)$ together with the associated normal orientations and the integers as above, and we call the pair $(N,\mathcal{D})$ (or simply, $\mathcal{D}$) a normally oriented indexed diagram (or a diagram, in short). For each point $c_{\ell}\in C$, we call its image $\gamma(c_{\ell})$ a cuspidal point of $\mathcal{D}$.
For example, if $f:M\to N$ is a stable map, then $(N,f(S(f)))$ together with the normal orientations and the absolute indices as described in the previous section can be considered to be a diagram.
In order to make the descriptions simpler, we adopt the convention that
  1. (1)
    for $S_{k}$, the associated integer $i_{k}$ can be chosen from the set $\{0,1,\ldots,n-1\}$,
  2. (2)
    for $c_{\ell}$, the associated integer $j_{\ell}$ can be chosen from the set $\{0,1,\ldots,n-2\}$,
  3. (3)
    if we reverse the normal orientation of $\gamma|_{S_{k}}$, then the index $i_{k}$ turns into $(n-1)-i_{k}$, and the indices $j_{\ell}$ for the adjacent cuspidal points turn into $(n-2)-j_{\ell}$: i.e., such two diagrams are considered to be the same.
Note that for a cusp point, we often omit the index and we do not associate a normal orientation. This is because the normal orientation and the index of a cuspidal point can be read-off from the normal orientations and the indices of the two adjacent branches of curves.
Furthermore, two normally oriented indexed diagrams are isotopic if there exists an ambient isotopy of $N$ which takes one to the other including the normal orientations and the indices. We often use the terminology “diagram” in order to refer to its isotopy class.
Definition 3.2.
A move for a normally oriented indexed diagram $(N,\mathcal{D})$ is an operation which changes $\mathcal{D}$ to a new diagram $\mathcal{D}^{\prime}$ such that
  1. (1)
    there exists a small disk $\Delta$ embedded in $N$ such that $\partial\Delta$ intersects $\mathcal{D}$ and $\mathcal{D}^{\prime}$ transversely outside of their cuspidal points and double points, and
  2. (2)
    $\mathcal{D}$ coincides with $\mathcal{D}^{\prime}$ outside of $\Delta$, including the normal orientations and the indices.
Since such an operation does not change anything outside of the disk $\Delta$, the general operation which replaces $\Delta\cap\mathcal{D}$ with $\Delta\cap\mathcal{D}^{\prime}$, is also called a move, and it is often represented by the symbol “$\Delta\cap\mathcal{D}\longrightarrow\Delta\cap\mathcal{D}^{\prime}$.”
Definition 3.3.
Suppose that a move $\Delta\cap\mathcal{D}\longrightarrow\Delta\cap\mathcal{D}^{\prime}$ is given. Let $f:M\to N$ be an arbitrary stable map such that
  • ()($\sharp$)( ♯ )
    there exists a disk $\Delta_{f}$ in $N$ such that for the diagram $(N,f(S(f)))$, $\partial\Delta_{f}$ intersects $f(S(f))$ outside of the images of the cusps and double points transversely and that $\Delta_{f}\cap f(S(f))$ is isotopic to $\Delta\cap\mathcal{D}$ under a certain identification of $\Delta_{f}$ with $\Delta$.
Suppose that there exists a smooth $1$–parameter family $f_{t}:M\to N$, $t\in[0,1]$, of smooth maps, with $f_{0}=f$, which modifies the singular point set image $f(S(f))$ only in $\mathop{\mathrm{Int}}\nolimits{\Delta_{f}}$, such that $\Delta_{f}\cap f_{1}(S(f_{1}))$ is isotopic to $\Delta\cap\mathcal{D}^{\prime}$, that $f_{t}^{-1}(\Delta_{f})=f^{-1}(\Delta_{f})$ for all $t$, that $f_{t}=f$ outside of the inverse images of $\Delta_{f}$ for all $t$, and that except for finitely many $t$’s, $f_{t}$ is a stable map. Then, we say that the move $\Delta\cap\mathcal{D}\longrightarrow\Delta\cap\mathcal{D}^{\prime}$ is realized for $f$ (or is realizable). If a move is realized for every $f$ with property ()($\sharp$)( ♯ ) above, then we say that the move is always-realizable.
In the following, we give several sets of moves that appear in a generic $1$–parameter family of smooth maps of closed $n$–dimensional manifolds, $n\geq 2$, into surfaces. Here, a smooth $1$–parameter family of smooth maps $\{f_{t}\}$ is said to be generic if every $f_{t}$ is a stable map except for a finite number of $t$’s. We usually impose a more technical condition that the family corresponds to a curve in the mapping space that is transverse to codimension $1$ unstable loci. For more technical details, see [3, 4, 5, 9, 21], for example.
Some of the following moves are always-realizable, while others are not. The reader should refer to the arguments given in [2], where the case of $n=4$ is thoroughly explained. The arguments for general dimension $n$ are basically similar, although there are some subtle issues for higher dimensions [8].
Fig. 4 gives the first set of moves that we will use later. Note that outside of the small disks depicted in the figure, these moves do not change the diagrams. These are called moves of type I: note that each move corresponds to a generic $1$–parameter family of smooth maps that changes the map only in a neighborhood of a point in the source manifold. The reader is referred to Chapter V and Chapter VII, Proposition 1 of [8] for $n\geq 6$.
[…]
Figure 4. Moves of type I
Note that the simple arrows are always-realizable, while the arrows with a slanted line are not. “Cusp Merge”, which is depicted by a dotted arrow, is “always-realizable” as long as the fibers over the region indicated by ()($\ast$)( ∗ ) of a given stable map are connected.
The moves as depicted in Fig. 5 are called moves of type II. Note that these moves correspond to generic $1$–parameter families of smooth maps that change the maps in neighborhoods of two points in the source manifold. For these moves and the following, the reader is referred to Chapter V and Chapter I, §7 of [8].
[…]
Figure 5. Moves of type II
The moves as depicted in Fig. 6 are called moves of type III. Note that these moves correspond to generic $1$–parameter families of smooth maps that change the maps in neighborhoods of three points in the source manifold.
[…]
Figure 6. Moves of type III
The moves as depicted in Fig. 7 are called cusp–fold crossing moves (refer to [8, Chapter V, (0.3)]).
[…]
Figure 7. Cusp–fold crossing moves
For our purpose, we also need some combination moves which are realized by a finite sequence of realizable moves. They are called an exchange move and a criss-cross move (see Fig. 8). Note that these moves are “always-realizable” as long as the fibers are connected over the regions indicated by ()($\ast$)( ∗ ) of a given stable map. Therefore, they are depicted by dotted arrows.
[…]
Figure 8. Exchange move and criss-cross move
An exchange move is realized as in Fig. 9. A criss-cross move is realized by combining a move of type II and an exchange move.
[…]
Figure 9. Moves for the exchange move
A more detailed discussion about the exchange move can be found in [2, Proposition 3.2] for $n=4$ and $i=1$. The reader is also referred to [8, Chapter V, §4].
4. Stable maps into $S^{2}$
In this section, we state and prove our first main result concerning stable maps into $S^{2}$. Let us start with the following definition.
Definition 4.1.
A generic map $f:M\to N$ of a closed manifold into a surface is said to be image simple if $f|_{S(f)}$ is a topological embedding.
Note that an image simple generic map is always stable as long as the source manifold is closed.
Remark 4.2.
Recall that a generic map $f:M\to N$ is simple if it has no cusps and each connected component of every fiber contains at most one singular point (see [12]). In the literature, some authors use the terminology “strongly simple” instead of “image simple”.
In this section, we prove the following.
Theorem 4.3.
Let $M$ be a closed connected $n$–dimensional manifold with $n\geq 3$. Then, every map $M\to S^{2}$ is homotopic to a stable map $f:M\to S^{2}$ with the following properties.
  1. (1)
    When $n$ is odd, $f$ has only folds of absolute index $(n-1)/2$ and has no cusps.
  2. (2)
    When $n$ is even, $f$ has only folds of absolute index $(n-2)/2$ and at most one cusp, whose absolute index is equal to $(n-2)/2$; furthermore, $f$ is image simple.
Remark 4.4.
The above theorem holds also for $n=2$. See [27, Theorem 1.4], for example.
Remark 4.5.
As the following proof shows, starting from an arbitrary stable map $M\to S^{2}$, we have an explicit algorithm, as a sequence of always-realizable moves, that modifies the given map to the one as described in the theorem.
For the moment, let us consider a stable map $f:M\to N$, where $M$ is a closed manifold of dimension $n\geq 3$ and $N=S^{2}$ or $\mathbb{R}^{2}$. In the following, for a closed manifold $M$, $\chi(M)$ denotes the Euler characteristic of $M$.
According to Thom [23], the number of cusp points of $f:M\to N$ has the same parity as $\chi(M)$. Levine [11] showed that if $M$ is connected, then every map $M\to N$ is homotopic to a generic map $f$ such that
  1. (i)
    $f$ has no cusp points when $\chi(M)$ is even, and
  2. (ii)
    $f$ has exactly one cusp point, whose index is equal to $(n-2)/2$, when $\chi(M)$ is odd (and in this case $n$ is necessarily even).
In fact, if we start with a stable map, then Levine showed that by applying always-realizable cusp-fold crossing moves and cusp merge moves, one can modify any given stable map homotopically to get one satisfying the above conditions (i) and (ii).
Therefore, for the proof of Theorem 4.3, we may assume that $f$ has at most one cusp of index $(n-2)/2$ when $n$ is even, and no cusp when $n$ is odd.
Let us now consider the elimination of definite folds. The author [16, 17] showed that for a closed manifold $M$ of dimension $n\geq 3$, every map $M\to S^{2}$ is homotopic to a stable map without definite fold singularities. In fact, the proof shows that any given stable map with at most one cusp point can be modified by applying some flips, cusp merges, moves of types II and III, and the moves as depicted in Fig. 10 with $i_{0}=0$ so that we get a stable map with at most one cusp point and without definite folds.
[…]
Figure 10. Turning a fold circle of index $i_{0}$ to one of index $i_{0}+1$, which is possible when $i_{0}\leq(n-3)/2$.
Note that, clearly, this is, in general, impossible for maps into $\mathbb{R}^{2}$.
Remark 4.6.
Let us consider the neighborhood of the index 00 fold locus corresponding to Fig. 10 with $i_{0}=0$. If it is orientable, then we see that the newly created fiber over a point in the innermost region is diffeomorphic to $S^{1}\times S^{n-3}$. If it is non-orientable, then we see that the fiber is diffeomorphic to the non-orientable $S^{n-3}$–bundle over $S^{1}$.
Then, for $n=3$, we see that the proof of Theorem 4.3 is complete. In the following, we assume $n\geq 4$.
Let us assume that $f:M\to S^{2}$ is a stable map with at most one cusp such that the minimum of the absolute indices of folds is equal to $i_{0}$ with $0<i_{0}<\lfloor(n-1)/2\rfloor$. We successively eliminate fold singularities (by homotopy) in the order of absolute index, without changing the number of cusps, as follows.
First, let us define the following notion.
Definition 4.7.
Let $f:M\to N$ be a stable map into a surface $N$. We say that the fold image of $f$ of absolute index $i$ is outward-directed (resp. inward-directed), if the image of all the folds of absolute index $i$ of $f$ is contained in the interior of a $2$–disk $D$ in $N$ in such a way that the complement of a regular value $z_{0}\in D$ can be nonsingularly foliated by arcs oriented from $z_{0}$ to $\partial D$, which intersect the image of each fold arc of absolute index $i$ transversely in its normal direction (resp. the opposite direction). In this case, the connected component of the complement of the fold image of absolute index $i$ in $D$ containing $z_{0}$ is called the innermost region.
Note that when $f$ has exactly one cusp point of index $(n-2)/2$, if $f$ is outward-directed, then the cusp is also “outward” (see Fig. 3 (3)).
We first show the following, which makes the fold image of absolute index $i_{0}$ outward-directed. We warn the reader that the case with $i_{0}=\lfloor(n-1)/2\rfloor$ is excluded.
Lemma 4.8.
Let $f:M\to S^{2}$ be a stable map with at most one cusp point such that the minimum of the absolute index of folds is equal to $i_{0}$ with $0<i_{0}<\lfloor(n-1)/2\rfloor$. Then, by using always-realizable moves we can make the fold image of absolute index $i_{0}$ outward-directed in such a way that the other folds have absolute index strictly greater than $i_{0}$, that their images are situated in the innermost region and that the resulting stable map has at most one cusp.
Proof.
Let $D\subset S^{2}$ be a $2$–disk containing $f(S(f))$. We can also take an annulus $A\cong[0,1]\times S^{1}$ in $D$ containing $f(S(f))$ in its interior. We use almost the same argument as in the proof of [2, Theorem 4.1] in order to arrange so that the fold image of absolute index $i_{0}$ is contained in $A$, and is outward-directed, and that the other fold images are situated in the inner region of $D\setminus A$. The points different from the situation in [2] that we need to take care of are as follows.
(1) In [2], the dimension of the source manifold $n$ was equal to $4$ and we had only folds of absolute index $i_{0}=1$. In our present case, the dimension is greater than or equal to $4$ and we may have folds of absolute index $>i_{0}$. When we apply always-realizable moves as in [2], as the folds other than those of absolute index $i_{0}$ have absolute index strictly greater than $i_{0}$, we can safely apply the necessary always-realizable moves.
(2) In [2], there were Lefschetz singular points and one could use such singular points in order to eliminate cusps, while in our situation we have no Lefschetz singular points, but may have a cusp point. If we have a cusp, then $n$ is even and the absolute index of the cusp must be equal to $(n-2)/2$, which is maximal. Therefore, we can use cusp-fold crossing moves that are always-realizable instead of always-realizable moves of type II.
(3) In [2], after applying a flip that creates two cusps, we eliminated the cusps by using Lefschetz singular points. In our situation, we can realize the same procedure as in Fig. 11 by creating a fold circle of absolute index $i_{0}+1$ whose image is embedded and is outward-directed. Note that the cusp merge can be performed, since there exists a small $2$–disk containing the two cusp images such that the two cusps lie on the same connected component of the inverse image of the $2$–disk.
[…]
Figure 11. Eliminating a self fold crossing
Details of the rest of the proof are left to the reader. ∎
Remark 4.9.
(1) In fact, we could perform the above procedure before eliminating the definite folds. In that case, we get a stable map as in the lemma with $i_{0}=0$.
(2) The above lemma is for maps into $S^{2}$: however, as the proof shows, the necessary modifications are performed only in the inverse images of the $2$–disk $D$. The map outside that part is unchanged. Therefore, for example, the regular fibers outside of $D$ stay the same.
(3) When $i_{0}=\lfloor(n-1)/2\rfloor$, the arguments in the proof does not work, since $i_{0}+1$ is no longer an absolute index. Furthermore, when $n$ is odd and $i_{0}=\lfloor(n-1)/2\rfloor$, we cannot apply type II moves as in the above argument, since reversing the normal orientation of the index $i_{0}$ fold image does not change the index.
Then, we need the following in order to eliminate folds of absolute index $i_{0}$.
Lemma 4.10.
Let $f:M\to S^{2}$ be a stable map with at most one cusp point such that the minimum of the absolute index of folds is equal to $i_{0}$ with $0<i_{0}<\lfloor(n-1)/2\rfloor$. We assume that $M$ is connected. Then, by using always-realizable moves we can eliminate the folds of absolute index $i_{0}$ in such a way that the other folds have absolute index strictly greater than $i_{0}$ and that the resulting stable map has at most one cusp.
Proof.
First, by Lemma 4.8, we may assume that the fold image of absolute index $i_{0}$ is outward-directed in such a way that the other folds have absolute index strictly greater than $i_{0}$, and that their images are situated in the innermost region, which is homeomorphic to an open disk and is denoted by $R$. Since $f$ has directed fold image of index $i_{0}$, we may assume that it winds around the north pole $NP\notin R$ of $S^{2}$ and that it is normally oriented toward $NP$.
Note that if there is a cusp, then the unique cusp lies on a component adjacent to fold loci whose absolute indices are strictly greater than $i_{0}$. In the following argument, we only modify the fold image of index $i_{0}$, so that we do not move the fold image adjacent to the cusp image.
Now, observe that the fibers over the points in $R$ are all connected. This can be seen as follows. As the absolute indices of folds whose images lie inside $R$ are greater than or equal to $i_{0}+1$, we see that when we cross a fold image, the topological change of fibers can be described by a surgery corresponding to the attachment of a handle of index $i$ with $32\leq i\leq n-$ 3. Therefore, the number of connected components of fibers does not change. If we consider the topological changes of fibers toward $NP$, then they correspond to the attachments of handles of index $n-1-i_{0}$, since the fold image of index $i_{0}$ is normally oriented toward $NP$. This means that the numbers of connected components of fibers may increase (if $i_{0}=1$); however, they never strictly decrease. Therefore, if a fiber over a point in $R$ is disconnected, then $f^{-1}(R)$ is disconnected, since $R$ is simply connected; hence, the manifold $M$ must necessarily be disconnected. This is a contradiction. Therefore, the fibers over the points in $R$ are all connected.
Then, by an argument similar to that in [2], we can arrange the fold image of index $i_{0}$ so that the image of each fold component of index $i_{0}$ winds exactly once around $NP$, by using moves of type II together with criss-cross moves.
We further follow the procedure as described in [2, Figs. 30 and 31]. If the top strand, denoted by $b$, does not intersect the other strands, then we can isotope it into a small open disk neighborhood of the north pole $NP$. Then, we can use the move as described in Fig. 10 of the present paper so as to turn it to a fold image of absolute index $i_{0}+1$ which is an embedded small circle and is directed outward. If the top strand does intersect the others, then we perform the moves as described in [2, Fig. 33], by using the move as in Fig. 11 of the present paper instead of the move as described in [2, Fig. 34]. Note that by this move, we again create an embedded fold image circle of index $i_{0}+1$, which can be isotoped to a neighborhood of the south pole $SP$. In order to reach the second stage of [2, Fig. 33], we need to swing the relevant fold image of index $i_{0}$ over the region around $SP$: this is possible, since all the fold images in this region have absolute index strictly greater than $i_{0}$.
Then, we reach the final stage of [2, Fig. 35]. Finally, we apply the move as in Fig. 10 of the present paper to turn the index $i_{0}$ circle to one of index $i_{0}+1$.
Repeating these procedures for the rest of the fold components of absolute index $i_{0}$, we can eliminate all the folds of absolute index $i_{0}$ so that all the folds of the resulting stable map have absolute index strictly greater than $i_{0}$ and that it has at most one cusp. This completes the proof. ∎
When the dimension of the source manifold is even, we can further arrange the images of the folds and cusps as follows.
Lemma 4.11.
Suppose that $n\geq 4$ is even. Let $f:M\to S^{2}$ be a stable map with at most one cusp point such that the absolute indices of all folds are equal to $i_{0}=\lfloor(n-1)/2\rfloor=(n-2)/2$. We assume that $M$ is connected. Then, by using always-realizable moves we can make the fold image topologically embedded in such a way that all the folds have absolute indices equal to $i_{0}$ and that the resulting stable map has at most one cusp.
Proof.
We follow the same argument as in the proof of Lemma 4.10. However, we need to be careful, since if there is a cusp, then we need to move the fold loci adjacent to the unique cusp as well.
In our present case, there may appear some small embedded fold image circles of index $i_{0}+1$ normally oriented outward (see Figure 11). Note that if we reverse the normal orientation, then the index turns into $i_{0}$, the absolute index, since $n$ is even and $i_{0}=(n-2)/2$. We used push moves under the presence of Lefschetz critical points (see [2, Fig. 2]). In our present case, we can instead use the moves as depicted in Fig. 12.
[…]
Figure 12. Moves replacing a push move
Then, we reach the situation as described in the leftmost figure of [2, Fig. 36], where the absolute indices are all equal to $i_{0}$ and each component is normally oriented inward. Furthermore, one of the components may contain a cusp image.
This completes the proof. ∎
Proof of Theorem 4.3.
When $n$ is odd, by applying Lemma 4.10 successively, we get a desired stable map into $S^{2}$.
When $n$ is even, by applying Lemma 4.10 successively, we get a stable map into $S^{2}$ with at most one cusp such that all the folds have absolute index $(n-2)/2$. Then, by applying Lemma 4.11, we get a desired stable map.
This completes the proof. ∎
Remark 4.12.
In the case of dimension $4$ in [2], we have allowed the Lefschetz singularities modeled on the map germ
$(z_{1},z_{2})\mapsto z_{1}^{2}+z_{2}^{2}$
at the origin through $C^{\infty}$ complex coordinates, and used always-realizable moves involving such Lefschetz singularities as well. In higher even dimensions, say $n=2m$, we also have a similar singularity modeled on
$(z_{1},z_{2},\ldots,z_{m})\mapsto z_{1}^{2}+z_{2}^{2}+\cdots+z_{m}^{2}.$
It would be an interesting problem to use this kind of singularities together with folds and cusps to obtain another proof to the above results.
5. Maps into $\mathbb{R}^{2}$
In this section, we show the following.
Theorem 5.1.
Let $M$ be a closed $n$–dimensional manifold with $n\geq 4$, where $n$ is even. Then, there exists an image simple generic map $f:M\to\mathbb{R}^{2}$ with at most one cusp point.
Remark 5.2.
The above theorem holds also for $n=2$; in fact, for each closed connected surface $M$, we can explicitly construct a generic map as described in the theorem. When $M$ is orientable, this is straightforward. For the non-orientable case, we can first construct an image simple generic map into $S^{2}$ by using [27, Theorem 1.4 (2)], for example, and then compose it with an appropriate generic map $S^{2}\to\mathbb{R}^{2}$ to get a desired generic map.
Proof of Theorem 5.1.
There exists a generic map $f:M\to S^{2}$ which is image simple by Theorem 4.3. Let $D_{S}\cup A\cup D_{N}=S^{2}$ be a decomposition of $S^{2}$ such that $D_{S}$ and $D_{N}$ are disjoint $2$–disks, that $A$ is an annulus, that they are attached along their boundaries, and that $\mathop{\mathrm{Int}}\nolimits{D_{N}}\supset f(S(f))$. For a regular value $q\in D_{S}$, set $F=f^{-1}(q)$, which is an $(n-2)$–dimensional closed manifold. Note that $f|_{f^{-1}(D_{S}\cup A)}$ is a trivial smooth $F$–bundle over $D_{S}\cup A$ by Ehresmann’s fibration theorem. In particular, we have $f^{-1}(A)\cong F\times[-1,1]\times S^{1}$.
Let $h:F\to[1,2]$ be a Morse function such that the critical values are all distinct. Define the Morse function $\tilde{h}:F\times[-1,1]\to[1,3]$ by $\tilde{h}(x,t)=h(x)\cos(\pi t/2)+1$, $(x,t)\in F\times[-1,1]$ (refer to [2, Fig. 40]). Then, we can construct a stable map $g:M\to\mathbb{R}^{2}$ in such a way that $g|_{f^{-1}(D_{N})}=i_{N}\circ f|_{f^{-1}(D_{N})}$, $g|_{f^{-1}(D_{S})}=i_{S}\circ f|_{f^{-1}(D_{S})}$, and $g|_{f^{-1}}(A)$ is the composition of a trivial $S^{1}$–family of Morse functions $\tilde{h}\times\mathrm{id}_{S^{1}}:(F\times[-1,1])\times S^{1}\to[1,3]\times S^{1}$ and an appropriate embedding $[1,3]\times S^{1}\to\mathbb{R}^{2}$, where $i_{N}:D_{N}\to S^{2}$ and $i_{S}:D_{S}\to S^{2}$ are appropriate embeddings (see Fig. 13). For details, refer to [2, Proof of Theorem 7.1].
[…]
Figure 13. Construction of a stable map into the plane
The resulting map is an image simple generic map with at most one cusp. This completes the proof. ∎
6. Relation to open book structures
In this section, we given an application of the above results in the odd dimensional case to open book structures.
Definition 6.1.
Let $M$ be a closed $n$–dimensional manifold. Suppose that $B$ is a closed submanifold of $M$ of dimension $n-2$ with trivial normal bundle, and let $\tau:N(B)\to B\times D^{2}$ be a trivialization of the normal disk bundle $N(B)$ of $B$ in $M$, where $N(B)$ is identified with a closed tubular neighborhood of $B$ in $M$. We assume that there exists a smooth $S^{1}$–bundle $\varphi:M\setminus B\to S^{1}$ such that $\varphi|_{N(B)\setminus B}$ coincides with the composition
[…]
where $pr_{2}$ is the projection to the second factor and $r$ is the radial projection. Such a structure of $M$ is called an open book structure. Note that for each $\theta\in S^{1}$, the closure $F_{\theta}$ of $\varphi^{-1}(\theta)$ in $M$ coincides with $\varphi^{-1}(\theta)\cup B$ and is a compact $(n-1)$–dimensional submanifold of $M$ with boundary $B$. The submanifold $B$ is called the binding and each $F_{\theta}$ is called a page of the open book structure.
For some historical account of open book structures, the author is referred to Quinn [14, §2].
Definition 6.2.
Let $M$ be a compact $n$–dimensional manifold with boundary. Suppose that $B$ is a proper $(n-2)$–dimensional compact submanifold of $M$ with trivial normal bundle. Here a submanifold $B$ is proper if $B\cap\partial M=\partial B$ and $B$ intersects $\partial M$ transversely. Let $\tau:N(B)\to B\times D^{2}$ be a trivialization of the normal disk bundle $N(B)$ of $B$ in $M$, where $N(B)$ is identified with a closed tubular neighborhood of $B$ in $M$ such that $N(B)\cap\partial M$ can be identified with the tubular neighborhood $N(\partial B)$ in $\partial M$. We assume that there exists a smooth $S^{1}$–bundle $\varphi:M\setminus B\to S^{1}$ such that $\varphi|_{N(B)\setminus B}$ coincides with the composition (6.1) as above and that $\varphi|_{\partial M\setminus\partial B}$ induces an open book structure with binding $\partial B$ on $\partial M$. Such a structure of $M$ is also called an open book structure; $B$ is called the binding and the closure of each fiber of $\varphi$ is called a page. Note that the restriction of the structure to $\partial M$ gives rise to an open book structure of $\partial M$. In this case, we say that the open book structure on the boundary extends through $M$.
As an application of our results, we give a new proof to the following result originally due to Quinn [14] (see also [10, 22, 25]).
Theorem 6.3.
Let $M$ be a compact connected $n$–dimensional manifold with $n\geq 5$ odd such that an open book structure on $\partial M$ is given. Then, it always extends through $M$.
Proof.
First, let us construct a certain generic map $f:M\to\mathbb{R}^{2}$ compatible with the open book structure on $\partial M$ such that $f(M)$ coincides with the unit disk $\Delta\subset\mathbb{R}^{2}$, as follows.
Let $bB\subset\partial M$ be the binding, $b\tau:N(bB)\to bB\times D^{2}$ the trivialization, and $b\varphi:\partial M\setminus bB\to S^{1}$ the fibration associated with the open book structure on $\partial M$. Let $f|_{N(bB)}:N(bB)\to\mathbb{R}^{2}$ be defined by the composition
[…]
where $pr_{2}$ is the projection to the second factor and $\iota:D^{2}\to\mathbb{R}^{2}$ is an embedding onto $\Delta$. We further define $f|_{\partial M\setminus\mathop{\mathrm{Int}}\nolimits{N(bB)}}:\partial M\setminus\mathop{\mathrm{Int}}\nolimits{N(bB)}\to\partial\Delta$ as the composition of $b\varphi|_{\partial M\setminus\mathop{\mathrm{Int}}\nolimits{N(bB)}}:\partial M\setminus\mathop{\mathrm{Int}}\nolimits{N(bB)}\to S^{1}$ and an appropriate diffeomorphism $S^{1}\to\partial\Delta$ in such a way that $f|_{\partial M}$ thus constructed gives a smooth map $\partial M\to\Delta\subset\mathbb{R}^{2}$. Here, strictly speaking, $M$ is considered to be a manifold with corners along $\partial N(bB)$. Note that this map has no singular points and that it can naturally be extended to a submersion, denoted by $f$, near $\partial M$ so that $f^{-1}(\partial\Delta)=\partial M\setminus\mathop{\mathrm{Int}}\nolimits{N(bB)}$.
As $\Delta$ is contractible, we can further extend the map $f$ over the entire manifold $M$ as a smooth map and then approximate it by a generic map, still denoted by $f$. We may assume that $f$ satisfies the conditions (1) and (2) of Definition 2.3. As it is a submersion near $\partial M$, we see that $S(f)$ is contained in $\mathop{\mathrm{Int}}\nolimits{M}$ and we may further assume that its image $f(S(f))$ sits inside $\mathop{\mathrm{Int}}\nolimits{\Delta}$ and that $f(M)=\Delta$.
Note that the number of cusps of $f$ is even, since the Euler characteristic of the pair $(M,F)$ vanishes, where $F$ is a fiber of $f$ over a point in $\partial\Delta$. Then, by the same procedure as in Levine’s cusp elimination theorem [11], we can eliminate the cusps of $f$ by pairs, since $M$ is connected. So, we may assume that $f$ has no cusps.
Now, we apply Lemma 4.8, taking into account the comments in Remark 4.9 (1) and (2), first to the definite fold image. As a result, we get a generic map $M\to\mathbb{R}^{2}$, denoted by $f$ again, such that $f(M)$ coincides with $\Delta$, the definite fold image is outward directed, and that the other fold images lie inside the interior of a $2$–disk, say $\Delta_{1}$, contained in the innermost region. Then, we apply Lemma 4.8 to $f$ restricted to $f^{-1}(\Delta_{1})$ to get a new generic map $f$ such that the fold image of absolute index $1$ is contained in $\mathop{\mathrm{Int}}\nolimits\Delta_{1}$, that it is outward directed, and that the other fold images lie inside the interior of a $2$–disk, say $\Delta_{2}$, contained in the innermost region. Repeating this procedure, we get a generic map $f:M\to\mathbb{R}^{2}$ such that $f(M)=\Delta$, that the fold images of absolute indices $i$ with $0\leq i<\lfloor(n-1)/2\rfloor$ are all directed outward, that their positions are consistent with the order of absolute indices (smaller indices in the outer position), and that the fold image of absolute index $i_{0}=\lfloor(n-1)/2\rfloor$ lies in the interior of a $2$–disk $\Delta_{i_{0}}$ contained in the innermost region. We may continue to assume that $f$ satisfies the conditions (1) and (2) of Definition 2.3.
Note that as $n$ is odd, even if we reverse the normal orientation of the fold image of absolute index $i_{0}$, the index does not change.
Now, we apply the always-realizable moves as depicted in Figs. 1417. We may assume that the disks $\Delta_{0},\Delta_{1},\ldots,\Delta_{i_{0}}$ are concentric, i.e., their centers all coincide with the origin $\mathbf{0}\in\mathbb{R}^{2}$. We denote by $\pi:\Delta\setminus\{\mathbf{0}\}\to S^{1}$ the radial projection. Let $A$ be the annulus $\Delta_{i_{0}}\setminus\mathop{\mathrm{Int}}\nolimits{\Delta_{i_{0}+1}}$, where $\Delta_{i_{0}+1}$ is a concentric disk in $\mathop{\mathrm{Int}}\nolimits{\Delta_{i_{0}}}$. We may assume that the fold image of absolute index $i_{0}$ lies in $\mathop{\mathrm{Int}}\nolimits{A}$ and that $\pi\circ f|_{S_{i_{0}}}:S_{i_{0}}\to S^{1}$ is an $S^{1}$–valued Morse function with distinct critical values, where $S_{i_{0}}$ is the set of folds of absolute index $i_{0}$ of $f$. Then, for each of the critical points, we apply the type II moves as depicted in Fig. 14 so that the images of the critical points by $f$ lie inside $\mathop{\mathrm{Int}}\nolimits{\Delta_{i_{0}+1}}$. (Before doing so, we arrange $f(S_{i_{0}})$ so that all the crossings are pushed into the regions in $A$ outside of those regions in which $f(S_{i_{0}})$ is depicted in the figure.) Note that in our case, only the top type II move in Fig. 5 is allowed and that the bottom type II move is not. Note also that as $n$ is odd, even if we reverse the normal orientation for a fold image of absolute index $i_{0}=(n-1)/2$, the absolute index does not change.
[…]
Figure 14. Pushing the local minima and maxima into $\mathop{\mathrm{Int}}\nolimits{\Delta_{i_{0}+1}}$.
We orient $S^{1}=\partial\Delta$ counter clockwise, for the moment, then with respect to that orientation, the numbers of local minima and maxima of the Morse function $\pi\circ f|_{S_{i_{0}}}$ are the same. Then, by further applying the type II moves, we may assume that the images of the critical points by $f$ lie in pairs, “face-to-face” as in the first figures of Figs. 15 and 16. Here, we can show that the fibers of $f$ over the points in $\Delta_{i_{0}+1}$ are connected by the same argument as in the proof of Lemma 4.10. Then, we apply always-realizable moves as depicted in Fig. 15 for each pair of the images of a local minimum and a local maximum. More precisely, we first apply type I moves to create 4 cusps. Then, we apply the cusp merge moves to the 2 pairs of cusps, which are allowed, since the fibers are connected. If we reverse the normal orientation of the newly born index–$(i_{0}+1)$ circle, then it turns into a fold image circle of absolute index $i_{0}-1$. Then, by applying the moves as depicted in Fig. 10 with $i_{0}$ replaced by $i_{0}-1$, we can turn this circle into a circle of absolute index $i_{0}$. Then, by applying the type II move once, we reach the final configuration in Fig. 15.
[…]
Figure 15. A series of moves for a pair of critical point images
By applying this procedure for each pair of critical point images, we reach the configuration as depicted in the figure on the right hand side in Fig. 16.
[…]
Figure 16. Applying the series of moves as in Fig. 15 for each pair
Then, inside $\mathop{\mathrm{Int}}\nolimits{\Delta_{i_{0}+1}}$, we move the circles using the type II moves several times as depicted in Fig. 17. We perform the moves in such a way that for the resulting map, still denoted by $f$, $\pi\circ f|_{S_{i_{0}}}:S_{i_{0}}\to\partial\Delta=S^{1}$ is a submersion.
[…]
Figure 17. Arranging the circles toward the center $\mathbf{0}$
Let $\Delta_{i_{0}+2}$ be a small closed $2$–disk centered at $\mathbf{0}$ disjoint from $f(S(f))$. Then, as $f$ restricted to $f^{-1}(\Delta_{i_{0}+2})$ and to $f^{-1}(\Delta_{i_{0}+2})\cap\partial M$ are submersions, we see that $f|_{f^{-1}(\Delta_{i_{0}+2})}:f^{-1}(\Delta_{i_{0}+2})\to\Delta_{i_{0}+2}$ is a trivial bundle. Furthermore, setting $B=f^{-1}(\mathbf{0})$, we see that $\partial B=bB$ and that the composition
[…]
is a smooth fiber bundle, since $\pi\circ f$ restricted to $S(f)$ is a submersion. Hence, this gives an open book structure for $M$. Making $N(bB)$ smaller, we see that an open book structure naturally isomorphic to the given one extends through $M$. This completes the proof. ∎
Remark 6.4.
Note that Theorem 6.3 also covers the case with $\partial M=\emptyset$: just ignore the boundary open book structures in the proof.
7. Elimination of indefinite folds of certain indices
In this section, using the constructions of the previous section for the odd dimensional case, we show that if $M$ is an odd dimensional closed manifold which is $k$–connected for a certain positive integer $k$, then we can eliminate folds of absolute indices $i$ with $1\leq i\leq k$ for maps into $\mathbb{R}^{2}$.
Let us start with the following.
Lemma 7.1.
For the extended open book structure on $M$ constructed in the above proof of Theorem 6.3, the binding $B$ is connected, and its fundamental group satisfies $\pi_{1}(M)\cong\pi_{1}(B)$ if $n\geq 7$.
Proof.
The binding $B$ is connected as shown in the proof of Theorem 6.3. Let us consider the fundamental group.
Take the segment $J=[0,1]\times\{0\}\subset\Delta$. We see that $F=f^{-1}(J)$ coincides with a page of the open book structure of $M$ and that $f|_{F}:F\to J=[0,1]$ is a Morse function such that $(f|_{F})^{-1}(0)=B=\partial F$. Furthermore, the indices of the critical points are all greater than or equal to $i_{0}=(n-1)/2\geq 3$. Hence, the inclusion map $B\to F$ induces an isomorphism between the fundamental groups.
Let us decompose $\Delta$, the unit disk in $\mathbb{R}^{2}$, into two half disks as $\Delta=\Delta_{+}\cup\Delta_{-}$, where $\Delta_{+}=\{(x,y)\in\Delta\,|\,y\geq 0\}$ and $\Delta_{-}=\{(x,y)\in\Delta\,|\,y\leq 0\}$. Then, we see that the inclusions $F\to f^{-1}(\Delta_{\pm})$ are homotopy equivalences. Now, $X=f^{-1}(\Delta_{+})\cap f^{-1}(\Delta_{-})$ is diffeomorphic to the union of $F$ and its copy attached along $B$ and hence the inclusion $B\to X$ induces an isomorphism between the fundamental groups by the van Kampen theorem. Finally, by the van Kampen theorem applied to $M=f^{-1}(\Delta_{+})\cup_{X}f^{-1}(\Delta_{-})$, we see that the inclusion $B\to M$ induces an isomorphism between the fundamental groups. This completes the proof. ∎
Theorem 7.2.
Let $M$ be a closed simply connected $n$–dimensional manifold with $n\geq 7$ odd. Then, there exists a stable map $f:M\to\mathbb{R}^{2}$ without cusp points such that the absolute indices $i$ of the folds are either equal to 00 or satisfy $22\leq i\leq(n-1)/$ 2.
Proof.
Let us consider the generic map $f:M\to\mathbb{R}^{2}$ constructed in the proof of Theorem 6.3. This is a generic map without cusp points; however, it may contain fold points of absolute index $1$. We will eliminate such fold points by homotopy as follows.
As in the proof of Lemma 7.1, we consider the Morse function $g=f|_{F}:F\to[0,1]$, where $F$ is a compact $(n-1)$–dimensional manifold with boundary $\partial F=B$. By the construction of $f$, we see that the indices $i$ of the critical points satisfy $(n-1)/2\leq i\leq n-1$. On the other hand, by our assumption, $B$ and $F$ are simply connected.
If $g$ has two or more critical points of index $n-1$, i.e. local maxima, then there exists a critical point of index $n-2$ in such a way that its dual handle connects two dual 00–handles, since $F$ is connected. Therefore, either of the two critical points of index $n-1$ and the index $n-2$ critical point make a canceling pair. (Note that by reversing the orientation of $J=[0,1]$, they make a canceling 001111 pair of critical points.) Then by [8, Part I, Chapter V, Proposition 1.1], we can cancel them leaving a pair of cusps of absolute index 00. By repeating this procedure, we may assume that $g$ has only one maximal point (refer to Fig. 18).
[…]
Figure 18. Canceling 001111 pairs of critical points
Let $U$ be a small open regular neighborhood of $J$ in $\mathbb{R}^{2}$. Note that $\Delta\setminus U$ is naturally identified with $[0,1]\times[-1,1]$, where the first factor corresponds to a radial outward arc and the second factor corresponds to the direction of $\partial\Delta=S^{1}$. Then, $f$ restricted to $f^{-1}(\Delta\setminus U)$ is considered to be a smooth $1$–parameter family of functions to $[0,1]$ parametrized by $[-1,1]$. Now, by applying an argument similar to that in the proof of [8, Part I, Chapter V, Proposition 3.5], we can homotopically modify $f$ so that it has only one fold circle of absolute index 00 and that it is embedded into $\mathbb{R}^{2}$ in the outermost place with the expense of creating some fold circles of absolute index $2$.
Since $F$ is simply connected of dimension greater than or equal to $6$, by [13, Theorem 8.1], we can cancel each critical point of index $1$ with the expense of creating critical points of index $3$. Then, again by the same argument as in the proof of [8, Part I, Chapter V, Proposition 3.5], we can homotopically modify $f$ so that it has no fold circle of absolute index $1$ and that it has only one fold circle of absolute index 00 embedded into $\mathbb{R}^{2}$ in the outermost place. This completes the proof. ∎
Lemma 7.3.
Let $M$ be a closed connected $n$–dimensional manifold, $n\geq 3$, and $f:M\to\mathbb{R}^{2}$ a generic map without fold of absolute index $1$. Then, the fundamental group $\pi_{1}(M)$ of $M$ is a free group of finite rank.
Proof.
Let us consider the quotient space $W_{f}$ as follows. For two points $x,x^{\prime}\in M$, we define $x\sim x^{\prime}$ if $f(x)=f(x^{\prime})$ and $x,x^{\prime}$ lie on the same connected component of $f^{-1}(f(x))=f^{-1}(f(x^{\prime}))$. Then $W_{f}$ is defined to be the quotient space $M/\!\sim$. Note that for the quotient map $q_{f}:M\to W_{f}$, there exists a unique continuous map $\bar{f}:W_{f}\to\mathbb{R}^{2}$ such that $f=\bar{f}\circ q_{f}$.
We may assume that $f$ is a stable map. Then, as $f$ does not have folds of absolute index $1$, we can show that $W_{f}$ has the structure of a compact connected smooth $2$–dimensional manifold (for example, see [26, Theorem 9.1.7] or [18]). Note that the boundary $\partial W_{f}$ coincides with the $q_{f}$–image of the definite fold locus. As we can see easily that $\bar{f}:W_{f}\to\mathbb{R}^{2}$ is an immersion, $W_{f}$ is orientable and has nonempty boundary. Let $\alpha_{1},\alpha_{2},\ldots,\alpha_{r}$ be disjoint properly embedded arcs in $W_{f}$ such that $Y=W_{f}\setminus\left(\cup_{s=1}^{r}\mathop{\mathrm{Int}}\nolimits{N(\alpha_{s})}\right)$ is diffeomorphic to the $2$–dimensional disk (with $4r$ corner points), where $N(\alpha_{s})$ are disjoint small regular neighborhoods of $\alpha_{s}$ in $W_{f}$. We may assume that the images of the cusp points by $q_{f}$ and the crossings of $q_{f}|_{S(f)}$ lie in $\mathop{\mathrm{Int}}\nolimits{Y}$, that $q_{f}|_{S(f)}$ is transverse to each $\alpha_{s}$, and that each connected component of $q_{f}(S(f))\cap N(\alpha_{s})$ is a properly embedded arc in $N(\alpha_{s})$ intersecting with $\alpha_{s}$ exactly at one point (see the top figure of Fig. 19).
[…]
Figure 19. Cutting open the quotient space
Let us cut open $W_{f}$ along $N(\alpha_{s})$, $s=1,2,\ldots,r$ so that we obtain the bottom figure of Fig. 19. Let $\beta$ be a properly embedded arc in $Y$ as in the figure. First, note that for each $s$, $M_{s}=q_{f}^{-1}(\alpha_{s})$ is a closed connected $(n-1)$–dimensional manifold and that $q_{f}|_{M_{s}}:M_{s}\to\alpha_{s}\cong[-1,1]$ is a Morse function with exactly one minima and one maxima and without critical points of index $1$ or $n-2$. Hence $M_{s}$ is simply connected. Let $h:Y\to\mathbb{R}$ be a smooth function which corresponds to the negative of the hight function as depicted in the bottom figure of Fig. 19. Then we may assume that $h\circ q_{f}:q_{f}^{-1}(Y)\to\mathbb{R}$ is a Morse function and it has exactly $r-1$ critical points of index $1$, each of which corresponds to a $1$–handle connecting $M_{s_{1}}$ and $M_{s_{2}}$, $s_{1}\neq s_{2}$. Hence, $q_{f}^{-1}(\beta)$ is diffeomorphic to the connected sum $M_{1}\sharp M_{2}\sharp\cdots\sharp M_{r}$, which is simply connected.
By considering the Morse function $h\circ q_{f}:q_{f}^{-1}(Y)\to\mathbb{R}$, we see that $q_{f}^{-1}(Y)$ is also simply connected. Thus, the original manifold $M$ is obtained from the simply connected manifold $q_{f}^{-1}(Y)$ by attaching simply connected manifolds $q_{f}^{-1}(N(\alpha_{s}))\cong M_{s}\times[-1,1]$ along their boundaries, $s=1,2,\ldots,r$. Hence, by an argument using the van Kampen theorem, we see that the fundamental group of $M$ is a free group of rank $r$. This completes the proof. ∎
Remark 7.4.
The above lemma implies that if we drop the condition that $M$ should be simply connected in Theorem 7.2, the conclusion does not hold in general: if $\pi_{1}(M)$ is not a free group, a generic map as described in the theorem does not exist. On the other hand, if $\pi_{1}(M)$ is a nontrivial free group, then we do not know if a generic map as described in the theorem exists or not.
Lemma 7.5.
For the extended open book structure on $M$ constructed in the proof of Theorem 6.3, if $M$ is $k$–connected for an integer $k$ with $21\leq k\leq(n-5)/$ 2, then the binding $B$ and the page $F$ are also $k$–connected.
Proof.
First note that $F$ is obtained by attaching handles of index $\geq(n-1)/2$ to $B\times[0,1]$, and hence that the inclusion map $B\to F$ induces isomorphisms in homology of dimensions $<(n-3)/2$.
Set $X_{\pm}=f^{-1}(\Delta_{\pm})$ and $X=f^{-1}(\Delta_{+})\cap f^{-1}(\Delta_{-})$. Then, we have the following Meyer–Vietoris exact sequence of homology with coefficients in $\mathbb{Z}$:
(7.1) $\cdots\to H_{i+1}(M)\to H_{i}(X)\to H_{i}(X_{+})\oplus H_{i}(X_{-})\to H_{i}(M)\to\cdots.$
We also have the exact sequence
$\cdots\to H_{i+1}(X)\to H_{i}(B)\to H_{i}(F)\oplus H_{i}(F)\to H_{i}(X)\to\cdots.$
As $H_{i}(B)\to H_{i}(F)\oplus H_{i}(F)$ is injective for $i<(n-3)/2$, we have the short exact sequence
$0\to H_{i}(B)\to H_{i}(F)\oplus H_{i}(F)\to H_{i}(X)\to 0$
This implies that the inclusion induces an isomorphism $H_{i}(F)\to H_{i}(X)$ for $i<(n-3)/2$. Thus, the inclusions $X\to X_{\pm}\simeq F$ also induce isomorphisms in homology. Hence, by (7.1) we have the short exact sequence
$0\to H_{i}(X)\to H_{i}(X_{+})\oplus H_{i}(X_{-})\to H_{i}(M)\to 0$
This implies that the inclusion induces an isomorphism $H_{i}(F)\to H_{i}(M)$. Since $B$ and $F$ are simply connected by Lemma 7.1 and $M$ is $k$–connected, we see that $B$ and $F$ are also $k$–connected. This completes the proof. ∎
Then, using the above lemma and an argument similar to that in the proof of Theorem 7.2, we get the following.
Theorem 7.6.
Let $M$ be a closed simply connected $n$–dimensional manifold with $n\geq 7$ odd. Suppose $M$ is $k$–connected for an integer $k$ with $21\leq k\leq(n-5)/$ 2. Then, there exists a stable map $f:M\to\mathbb{R}^{2}$ without cusp points such that $f$ has no fold points of absolute index $i$ with $1\leq i\leq k$.
Remark 7.7.
The author [15] showed that a closed orientable $3$–dimensional manifold $M$ admits an image simple generic map $f:M\to N$ without cusp points for a surface $N$ if and only if $M$ is a graph manifold. Note that a closed orientable $3$–dimensional manifold is a graph manifold if it is the union of some $S^{1}$–bundles over compact surfaces attached along the torus boundaries. Note that the class of graph manifolds is relatively small in the set of closed orientable $3$–dimensional manifolds: for example, hyperbolic $3$–manifolds never admit such an image simple generic map.
Furthermore, by [19], for every “generic map” $f:M\to N$ of a closed oriented $4$–dimensional manifold $M$ into an orientable $3$–dimensional manifold $N$, the restriction $f|_{S(f)}$ has at least $|\sigma(M)|$ triple points, where $\sigma(M)\in\mathbb{Z}$ denotes the signature of $M$.
Furthermore, Gromov [7] showed that for every $n\geq 3$, there exist a lot of $n$–dimensional closed manifolds $M$ such that for any generic map $f:M\to\mathbb{R}^{n-1}$, $f|_{S(f)}$ cannot be a topological embedding.
These results imply that for these dimensions, for many of the source manifolds, no map can be modified in such a way that the restriction to the singular point set is a topological embedding.
However, according to our results of this paper, when $n\geq 4$ is even, a map of a closed manifold of dimension $n$ into $S^{2}$ or $\mathbb{R}^{2}$ can always be homotopically modified so that it is an embedding on its singular point set.
Acknowledgment
The author would like to thank Professor İnanç Baykur for stimulating discussions which motivated the theme of this paper. He would also like to thank Dominik Wrazidlo and Rustam Sadykov for some discussions. The author would also like to thank Professor Vincent Blanlœil and the Advanced Mathematical Research Institute of the University of Strasbourg for the hospitality during the preparation of the manuscript. This work has been supported in part by JSPS KAKENHI Grant Numbers JP22K18267, JP23H05437. This work was also supported by the Research Institute for Mathematical Sciences, an International Joint Usage/Research Center located in Kyoto University.